# Beyond 1D: Vision Transformers and Multichannel Signal Images for PPG-to-ECG Reconstruction

### v1

@@ 1 Introduction
Electrocardiograms (ECGs) are essential tools for diagnosing and monitoring cardiovascular health, providing crucial insights into heart rate variability (HRV), heart rate, and key waveform features. These include the QRS complex, PR interval, ST segment, TP interval, and QT interval, which are vital for understanding the heart’s electrical activity and diagnosing various cardiac conditions [1, 2].
@@
For example, a normal PR interval ranges from 120 to 200 milliseconds. Prolonged PR intervals may indicate first-degree atrioventricular (AV) block or delayed conduction through the AV node, suggesting potential cardiac conduction issues [3]. Conversely, shortened PR intervals might imply conditions such as Wolff-Parkinson-White (WPW) syndrome or Lown-Ganong-Levine syndrome, where accessory pathways bypass the normal AV nodal delay [4].
@@
Similarly, prolonged ST segments are indicative of myocardial ischemia or acute myocardial infarction, resulting from delayed repolarization [5], or may reflect ventricular conduction defects that affect the timing of repolarization [5]. On the other hand, abnormally short ST segments can be a sign of hyperkalemia, which typically shortens ventricular repolarization [6], or congenital short QT syndrome, a rare but serious condition associated with a high risk of arrhythmias [7]. Recently, the ECG signal has been used for noninvasive serum electrolyte prediction and monitoring [8, 9].
@@
Despite the clinical significance of detailed waveform features, current research on ECG reconstruction from photoplethysmography (PPG) primarily concentrates on HRV and heart rate. These studies often struggle to accurately reconstruct smaller, yet critical, waveform features such as T and P waves. To bridge this gap, we introduce a novel approach focused on the precise reconstruction of these ECG waveform features. Our method is designed to facilitate a more comprehensive evaluation and improve diagnostic capabilities by prioritizing the accurate prediction and measurement of these essential waveforms.
@@
Reconstructing ECG signals from PPG addresses several limitations of traditional ECG monitoring, such as restricted mobility, skin irritation, and the need for offline data processing [10]. PPG is a non-invasive, wearable-friendly technology that provides valuable cardiovascular insights, making it ideal for long-term, real-time monitoring [11, 12, 13, 14]. Despite its benefits, ECG remains the gold standard for cardiovascular diagnostics due to its unmatched ability to detail critical waveform features necessary for clinical analysis [15]. The intrinsic physiological relationship between ECG and PPG signals, governed by the heart’s electrical activity and its influence on peripheral blood volume, serves as the foundation for reconstructing ECG signals from PPG [16, 13, 17].
Mapping the complex, non-linear relationship between ECG and PPG signals requires advanced signal processing and machine learning techniques. This task is particularly challenging due to signal artifacts and the inherent complexity of the ECG-PPG relationship. Although recent advancements in deep learning and hybrid models have shown promising results, most existing studies primarily rely on single-channel PPG signals [18, 10, 19, 20, 21, 15, 22, 23, 24, 25], which limits the diversity and richness of extracted features. In response, we propose a novel approach that utilizes multi-channel 2D PPG images for ECG reconstruction. The PPG image is constructed by stacking multiple padded PPG cycles into a 2D image representation. This representation consists of four channels: the original PPG signal image, its first-order difference image, second-order difference image, and area under the curve (AUC) image. This multi-channel design enables our model to learn more informative PPG representations and more effectively capture subtle waveform features.
Traditional convolutional neural networks (CNNs) have demonstrated strong performance in signal processing tasks, including ECG reconstruction. However, CNNs are inherently constrained by their reliance on localized operations and limited receptive fields, which restrict their ability to capture long-range dependencies and global relationships in the data. In contrast, recent advancements in self-attention-based models, such as Transformers [26], overcome these limitations by leveraging global context and effectively modeling sequential dependencies. Vision Transformers (ViTs), which adapt the Transformer architecture for image-based tasks, have achieved outstanding performance in analyzing spatially structured data. Building on these advancements, we leverage ViT for ECG reconstruction by applying it to our PPG signal images. This allows the ViT model to capture variations in PPG waveforms across multiple cycles within the signal images. Unlike CNN-based approaches, which focus on local features, ViTs enable the model to learn dependencies through self-attention mechanisms across both the temporal dimension (capturing sequential dependencies along the beat progression) and the inter-beat dimension (modeling relationships across multiple aligned beats in the 2D representation).
In this study, we propose a novel ECG reconstruction method based on a ViT network that processes four-channel signal images derived from PPG signals. The input data is represented as a 3D tensor with dimensions $X\times Y\times Z$, where $X$ and $Y$ correspond to a 2D signal image derived from the 1D PPG sequence, constructed by stacking padded PPG cycles (or beats) into a structured 2D representation, and $Z$ represents the number of channels.
Transforming the 1D sequence into a 2D representation enables the application of 2D patching in ViT, enhancing the model’s ability to capture structured temporal dependencies, where each patch retains both local waveform characteristics and global beat-to-beat variations. Using 2D patches is equivalent to applying a strided patching strategy in the original 1D sequence. This multi-dimensional representation allows ViT to capture both local and global temporal dependencies within each channel (via the X-Y plane) while simultaneously modeling inter-channel relationships (via the Z-axis). These additional channels, representing the first-order difference, second-order difference, and AUC of the PPG signal, enhance feature extraction by highlighting rate of change, signal curvature, and cumulative trends, which are critical for reconstructing ECG waveforms. By effectively learning complex dependencies across time and channels, our approach addresses the limitations of traditional 1D sequence-based models.
@@
    We introduce a novel four-channel signal image representation for ECG reconstruction, diverging from traditional methods that typically rely on single-channel PPG signals. Our model incorporates the original PPG signal, its first-order difference, second-order difference, and AUC, enriching the input data. This multi-channel approach allows the model to extract richer features, significantly enhancing the accuracy of ECG waveform predictions.
@@
    By integrating the PPG signal with its first-order difference, second-order difference, and AUC, our method captures more informative and physiologically relevant features essential for ECG reconstruction. Unlike previous PPG-to-ECG approaches, this multi-channel representation provides deeper insights into the intrinsic relationships between PPG and ECG signals.
@@
    We demonstrate the effectiveness of the ViT in ECG reconstruction, emphasizing its capability to process 2D signal representations. By capturing both temporal dependencies (across beats) and structured intra-beat variations, ViT effectively models complex relationships within and between beats. This enables it to extract richer PPG features and beat-to-beat variations for PPG-to-ECG mapping, achieving superior robustness and performance compared to recent state-of-the-art 1D sequence-based models.
@@
    We have developed new evaluation metrics tailored for a more detailed and clinically relevant assessment of ECG waveform reconstruction. These metrics include QRS area error, PR interval error, RT interval error, and RT amplitude difference, going beyond traditional measures like HRV and root mean square error (RMSE). These enhancements provide a more comprehensive and clinically aligned evaluation of waveform reconstruction quality.
@@ 2 Related Work
This section reviews recent advancements in methods for ECG reconstruction from PPG signals using 1D sequence-based approaches and introduces the Vision Transformer (ViT), which serves as the foundation of our proposed method.
@@ 2.2 Vision Transformer
In ViT, an input image with dimensions $H\times W\times C$ (height, width, channels) is divided into a grid of non-overlapping patches. Each patch consists of $P\times P$ pixels, where $P$ is a fixed patch size (e.g., $8\times 8$). The total number of patches, $N$, is computed as $N=\frac{H}{P}\times\frac{W}{P}$. Each patch is flattened into a 1D vector of size $P^{2}\times C$ by concatenating its pixel values, treating it as an individual token.
@@
with each attention head defined as:
$\text{head}_{i}=\text{Attention}(\mathbf{Q}\mathbf{W}_{i}^{Q},\mathbf{K}\mathbf{W}_{i}^{K},\mathbf{V}\mathbf{W}_{i}^{V}),\quad i=1,\dots,h,$
where the projection matrices are $\mathbf{W}_{i}^{Q},\mathbf{W}_{i}^{K},\mathbf{W}_{i}^{V}\in\mathbb{R}^{D\times d_{k}}$. Here, $d_{k}=d_{v}=D/h$, and $h$ is the number of attention heads. After computing the individual heads, the outputs are concatenated along the last dimension, yielding a tensor of shape $\mathbb{R}^{B\times T\times D}$. This concatenated output is then projected back to the original embedding dimension using an output projection matrix $\mathbf{W}^{O}\in\mathbb{R}^{D\times D}$.
Residual connections and layer normalization are subsequently applied to stabilize training and improve optimization:
$\mathbf{z}_{\text{att}}=\text{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})+\mathbf{z}_{\text{in}},$
where $\mathbf{z}_{\text{in}}$ denotes the input token embeddings to the transformer layer, and $\mathbf{z}_{\text{att}}$ is the output after applying multi-head self-attention.
To further refine the representation, the attention output is passed through a feed-forward network (FFN):
$\mathbf{z}_{\text{out}}=\text{FFN}(\mathbf{z}_{\text{att}}),$
where the FFN consists of two fully connected layers with a non-linear activation function:
$\text{FFN}(\mathbf{z}_{\text{att}})=\text{ReLU}(\mathbf{z}_{\text{att}}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2},$
Here, $\mathbf{z}_{\text{att}}\in\mathbb{R}^{B\times T\times D}$ denotes the output of the attention sub-layer. The weight matrix of the first linear layer is $\mathbf{W}_{1}\in\mathbb{R}^{D\times H}$, and its corresponding bias is $\mathbf{b}_{1}\in\mathbb{R}^{H}$, where $H$ is the hidden dimension of the feed-forward network. The second linear layer uses weight matrix $\mathbf{W}_{2}\in\mathbb{R}^{H\times D}$ and bias $\mathbf{b}_{2}\in\mathbb{R}^{D}$.
By treating image patches as tokens, ViT processes images in a manner similar to how transformers handle words in a sentence, allowing it to capture both local and global dependencies. Unlike CNNs, which inherently encode inductive biases such as locality and translation invariance, ViT learns these features directly from data. Its self-attention mechanism enables the modeling of long-range dependencies between distant patches, providing a more holistic understanding of global structures within an image, an aspect that CNNs may struggle to represent. This capability makes ViT particularly well-suited for ECG reconstruction, where accurately capturing both fine-grained waveform details and high-level beat patterns is crucial.
@@ 5.3 Experimental Results
We evaluated 30 record pairs from each dataset: BIDMC and CapnoBase, using a leave-one-out validation strategy, where one record served as the test set and the remaining were used for training. Because our evaluation metrics, including QRS area error, PR interval error, RT interval error, and RT amplitude difference, depend on peak-detection algorithms, and small waveforms (such as P waves) are particularly susceptible to noise, we selected 30 relatively low-noise recordings from each dataset to ensure reliable measurement accuracy. We used the state-of-the-art 1D convolutional model, CLEP-GAN [30], as our baseline. Table 1 summarizes the average performance on both datasets. All test results for our proposed method were based on the single-channel output structure shown in Figure 1.
@@
Figure 5 provides a visual comparison of the reconstructed ECG samples generated by our method and CLEP-GAN. In both examples, our method achieves markedly improved reconstruction, yielding signals that more closely resemble the original compared to those produced by CLEP-GAN. Additional reconstruction examples are provided in A.
@@ 6.0.1 Comparison of 2D Patch-Based and 1D Patch-Based Representations
We conducted a comparative analysis to assess the conditions under which the ViT with 2D patching outperforms the 1D patch-based Transformer in ECG reconstruction. Although both models share the same encoder-decoder architecture, they differ in their input representations. The ViT processes four-channel 2D patches, whereas the Transformer operates on four-channel 1D patches. In this configuration, the 1D sequences are segmented into sequential patches along the time axis.
Figure 7 illustrates the first eight patches extracted from a 2D PPG image and a 1D PPG signal, as used by the ViT and the Transformer networks, respectively. For better visualization, only the first channel of the image or signal is shown. In a 2D PPG image, each row represents a padded beat. When the patch size is set to $8\times 16$, each patch encompasses 16 consecutive points across 8 beats. Having multiple beats within each patch allows the ViT to effectively capture variations within similar segments in different cardiac cycles (as shown in Figure 7(b)). In contrast, with 1D patching (as shown in Figure 7(a)), each patch corresponds to a single segment from one beat, which limits the model’s capacity to capture inter-beat variations.
@@
As shown in Figure 10, the patches are flattened and sequentially arranged before being projected, along with position embeddings, into the transformer encoder. In our approach, the patch size is set to half the image height (i.e., the number of rows). Theoretically, larger patches contain richer information as they span multiple beats. However, using a patch size smaller than the total number of beats in an image allows the model to capture not only intra-beat but inter-beat relationships. Additionally, we observed that larger patch sizes reduce the total number of patches, leading to less stable outcomes.
Figure 9 compares ECG reconstructions generated by our proposed ViT-based method and the 1D patch-based Transformer. The results demonstrate that our method produces ECG signals significantly closer to the ground truth. This improvement is attributed to the ability of 2D patches to capture variations in ECG cycles, an essential capability that the 1D sequence-based Transformer lacks.

### v2

# Beyond Single-Channel: Multichannel Signal Imaging for PPG-to-ECG Reconstruction with Vision Transformers

@@ 1 Introduction
Electrocardiograms (ECGs) are essential tools for diagnosing and monitoring cardiovascular health, providing crucial insights into heart rate variability (HRV), heart rate, and key waveform features. These include the QRS complex, PR interval, ST segment, TP interval, and QT interval, which are vital for understanding the heart’s electrical activity and diagnosing various cardiac conditions [1, 2]. For example, a normal PR interval ranges from 120 to 200 milliseconds. Prolonged PR intervals may indicate first-degree atrioventricular (AV) block or delayed conduction through the AV node, suggesting potential cardiac conduction issues [3]. Conversely, shortened PR intervals might imply conditions such as Wolff-Parkinson-White (WPW) syndrome or Lown-Ganong-Levine syndrome, where accessory pathways bypass the normal AV nodal delay [3].
@@
Similarly, prolonged ST segments are indicative of myocardial ischemia or acute myocardial infarction, resulting from delayed repolarization, or may reflect ventricular conduction defects that affect the timing of repolarization [2]. On the other hand, abnormally short ST segments can be a sign of hyperkalemia, which typically shortens ventricular repolarization, or congenital short QT syndrome, a rare but serious condition associated with a high risk of arrhythmias [3]. Building on these diagnostic capabilities, recent research has explored using ECG signals for noninvasive serum electrolyte prediction and monitoring [4, 5].
@@
Despite the clinical significance of detailed waveform features, much of the existing research on ECG reconstruction from photoplethysmography (PPG) has primarily focused on estimating HRV and heart rate [6, 7]. However, these approaches often struggle to accurately recover smaller, yet diagnostically important, waveform components such as the T and P waves [8]. To address this limitation, we propose a novel method that emphasizes the precise reconstruction of these critical ECG features. By prioritizing accurate prediction and measurement of complete waveforms, our approach aims to enable more comprehensive evaluation and enhanced diagnostic capability.
@@
Reconstructing ECG signals from PPG addresses several limitations of traditional ECG monitoring, including restricted mobility, skin irritation, and the reliance on offline data processing [9]. As a non-invasive, wearable-compatible technology, PPG offers significant advantages for long-term and real-time cardiovascular monitoring [10]. For instance, Gil et al. [11] demonstrated the value of PPG in assessing vascular and respiratory functions, while Chua et al. [12] explored its application in nocturnal blood pressure estimation. Nevertheless, ECG remains the clinical gold standard due to its superior resolution of key waveform components crucial for cardiac diagnostics. The intrinsic physiological relationship between ECG and PPG, where cardiac electrical activity governs peripheral blood volume changes [13], underpins the rationale for PPG-to-ECG reconstruction models.
Drawing from established approaches in ECG analysis, the task of mapping the complex, non-linear relationship between ECG and PPG signals can be framed as a pattern recognition problem, with the objective of learning temporal and morphological correspondences between the two modalities [2]. This task is challenging due to signal artifacts and the inherent physiological differences: ECG captures the heart’s electrical activity, while PPG measures peripheral blood volume changes. Recent advances in pattern recognition, especially deep learning for time-series and physiological signal analysis, have shown significant promise in addressing these challenges [14, 15].
A variety of deep learning models have been proposed for PPG-to-ECG translation. For example, Zhu et al. [16] developed a convolutional neural network (CNN) to reconstruct ECG waveforms from single-channel PPG. Tang et al. [17] introduced an alignment-based LSTM model emphasizing temporal synchronization between PPG and ECG cycles. Tian et al. [9] proposed a joint dictionary learning framework to model cross-domain correspondences. Chiu et al. [18] employed attentional neural networks to reconstruct QRS complexes from PPG input. More recently, Shome et al. [19] proposed a region-disentangled diffusion model to selectively enhance high-information ECG segments, such as the QRS complex.
However, most existing methods rely on single-channel PPG signals, which limits their capacity to capture waveform variability across cardiac cycles. To overcome this limitation, we propose a multi-channel pattern recognition approach using 2D signal representations. Specifically, we generate a four-channel PPG image by stacking padded PPG cycles and computing three derived features: the first-order difference, the second-order difference, and the area under the curve (AUC). This enriched representation allows the model to capture both local waveform structures and inter-beat variability, thereby improving the accuracy and fidelity of ECG signal reconstruction.
Traditional CNNs have demonstrated strong performance in signal processing tasks, including ECG reconstruction. However, their reliance on localized filters and limited receptive fields constrains their capacity to capture long-range dependencies and global contextual patterns. Recent advances in self-attention-based models, particularly Transformers [20], have addressed these limitations by effectively modeling sequential dependencies through global context. These models have shown strong performance across a wide range of time-series applications [21, 22]. Vision Transformers (ViTs) [23], an extension of the Transformer architecture for image-based analysis, further enhance this capability by capturing spatial structures in data-rich image representations.
Building on these advancements, we propose a novel ECG reconstruction method that applies a ViT architecture to four-channel image-like representations of PPG signals. Unlike conventional 1D CNN-based models, which operate on raw PPG sequences, our method restructures the input as a 3D tensor of size $X\times Y\times Z$: the $X$-$Y$ plane captures a 2D signal image formed by stacking individually padded PPG cycles along temporal and morphological dimensions, while $Z$ denotes the number of feature channels. These include the raw PPG signal, its first-order difference, second-order difference, and the AUC.
This transformation from a 1D sequence to a structured 2D image enables the use of 2D patching in ViT, where each patch captures both local waveform morphology and cross-cycle dynamics. This 2D formulation implicitly incorporates strided segmentation along the temporal axis, similar to applying overlapping or non-overlapping patches in 1D sequence models. The $Z$-dimension further enriches this representation by incorporating derivative and cumulative signal features, improving the model’s ability to extract clinically relevant information. Through self-attention, the ViT captures dependencies within and across beats, modeling both intra-beat patterns and inter-beat temporal relationships. As a result, our approach addresses the key limitations of traditional 1D sequence-based models and offers enhanced fidelity in ECG waveform reconstruction.
@@
    We propose a novel four-channel image-based representation for ECG reconstruction, diverging from traditional single-channel PPG inputs. Our formulation combines the raw PPG signal with its first-order difference, second-order difference, and AUC, enriching the signal representation and enabling more comprehensive feature extraction.
@@
    By incorporating derivative and cumulative features of the PPG signal, our method captures physiologically meaningful dynamics that are critical for accurate ECG reconstruction. This multi-channel representation offers deeper insights into the temporal and morphological relationships between PPG and ECG signals.
@@
    We demonstrate the efficacy of ViTs for ECG reconstruction from structured 2D PPG representations. The ViT effectively captures intra-beat morphology and inter-beat temporal dependencies via self-attention, outperforming state-of-the-art 1D sequence-based models in robustness and accuracy.
@@
    We introduce new evaluation metrics: QRS area error, PR interval error, RT interval error, and RT amplitude difference, to assess ECG reconstruction quality beyond conventional metrics such as RMSE and HRV. These metrics provide a more nuanced and clinically interpretable assessment of waveform fidelity.
@@ 2 Related Work
This section reviews recent advancements in methods for ECG reconstruction from PPG signals using 1D sequence-based approaches and introduces the Vision Transformer, which serves as the foundation of our proposed method.
@@ 2.2 Vision Transformer
In ViT, an input image with dimensions $H\times W\times C$ (height, width, channels) is divided into a grid of non-overlapping patches. Each patch consists of $P\times P$ pixels, where $P$ is a fixed patch size (e.g., $8\times 8$). The total number of patches, $N$, is computed as $N=\frac{H}{P}\times\frac{W}{P}$.Each patch is then flattened into a 1D vector of size $P^{2}\times C$ by concatenating its pixel values, effectively treating each patch as a token.
@@
with each attention head defined as $\text{head}_{i}=\text{Attention}(\mathbf{Q}\mathbf{W}_{i}^{Q},\mathbf{K}\mathbf{W}_{i}^{K},\mathbf{V}\mathbf{W}_{i}^{V}),\quad i=1,\dots,h,$ where the projection matrices are $\mathbf{W}_{i}^{Q},\mathbf{W}_{i}^{K},\mathbf{W}_{i}^{V}\in\mathbb{R}^{D\times d_{k}}$. Here, $d_{k}=d_{v}=D/h$, and $h$ is the number of attention heads. After computing the individual heads, the outputs are concatenated along the last dimension, yielding a tensor of shape $\mathbb{R}^{B\times T\times D}$. This concatenated output is then projected back to the original embedding dimension using an output projection matrix $\mathbf{W}^{O}\in\mathbb{R}^{D\times D}$.
To stabilize training and facilitate optimization, residual connections and layer normalization are applied following the self-attention operation. The output of the multi-head attention layer is given by $\mathbf{z}_{\text{att}}=\text{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})+\mathbf{z}_{\text{in}},$ where $\mathbf{z}_{\text{in}}$ denotes the input token embeddings to the transformer layer, and $\mathbf{z}_{\text{att}}$ is the output after applying multi-head self-attention.
To further refine the representation, the attention output is passed through a feed-forward network (FFN): $\mathbf{z}_{\text{out}}=\text{FFN}(\mathbf{z}_{\text{att}}),$ where the FFN consists of two fully connected layers with a non-linear activation function, defined as $\text{FFN}(\mathbf{z}_{\text{att}})=\text{ReLU}(\mathbf{z}_{\text{att}}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2}.$ Here, $\mathbf{z}_{\text{att}}\in\mathbb{R}^{B\times T\times D}$ denotes the output of the attention sub-layer. The weight matrix of the first linear layer is $\mathbf{W}_{1}\in\mathbb{R}^{D\times H}$, and its corresponding bias is $\mathbf{b}_{1}\in\mathbb{R}^{H}$, where $H$ is the hidden dimension of the feed-forward network. The second linear layer uses weight matrix $\mathbf{W}_{2}\in\mathbb{R}^{H\times D}$ and bias $\mathbf{b}_{2}\in\mathbb{R}^{D}$.
By treating image patches as tokens, ViT processes images similarly to how transformers process sequences of words, allowing it to capture both local and global dependencies. Unlike CNNs, which incorporate strong inductive biases such as locality and translation equivariance, ViT learns such features purely from data. Its self-attention mechanism is particularly effective at modeling long-range dependencies between distant patches, allowing for a more holistic understanding of global image structure, an aspect that CNNs may struggle to represent. This capability makes ViT especially well-suited for ECG reconstruction, where both fine-grained waveform morphology and global rhythm patterns must be accurately captured.
@@ 5.3 Experimental Results
We evaluated 30 record pairs from each dataset: BIDMC and CapnoBase, using a leave-one-out validation strategy, where one record served as the test set and the remaining were used for training. Because our evaluation metrics, including QRS area error, PR interval error, RT interval error, and RT amplitude difference, depend on peak-detection algorithms, and small waveforms (such as P waves) are particularly susceptible to noise, we selected 30 relatively low-noise recordings from each dataset to ensure reliable measurement accuracy. As a baseline, we employed the state-of-the-art 1D convolutional model CLEP-GAN [28], chosen for its superior performance compared to other advanced models. Table 1 summarizes the average performance on both datasets. All test results for our proposed method were based on the single-channel output structure shown in Figure 1.
@@
Figure 5 provides a visual comparison of the reconstructed ECG samples generated by our method and CLEP-GAN. In both examples, our method achieves markedly improved reconstruction, yielding signals that more closely resemble the original compared to those produced by CLEP-GAN. Additional reconstruction examples are provided in B.
@@ 6.0.1 Comparison of 2D Patch-Based and 1D Patch-Based Representations
We conducted a comparative analysis to assess the conditions under which the ViT with 2D patching outperforms the 1D patch-based Transformer in ECG reconstruction. Although both models share the same encoder-decoder architecture, they differ in their input representations. The ViT processes four-channel 2D patches, whereas the Transformer operates on four-channel 1D patches. In this configuration, the 1D sequences are segmented into sequential patches along the time axis. Figure 7 illustrates the first eight patches extracted from a 2D PPG image and a 1D PPG signal, as used by the ViT and the Transformer networks, respectively. For better visualization, only the first channel of the image or signal is shown. In a 2D PPG image, each row represents a padded beat. When the patch size is set to $8\times 16$, each patch encompasses 16 consecutive points across 8 beats. Having multiple beats within each patch allows the ViT to effectively capture variations within similar segments in different cardiac cycles (as shown in Figure 7(b)). In contrast, with 1D patching (as shown in Figure 7(a)), each patch corresponds to a single segment from one beat, which limits the model’s capacity to capture inter-beat variations.
@@
As shown in Figure 10, the patches are flattened and sequentially arranged before being projected, along with position embeddings, into the transformer encoder. In our approach, the patch size is set to half the image height (i.e., the number of rows). Theoretically, larger patches contain richer information as they span multiple beats. However, using a patch size smaller than the total number of beats in an image allows the model to capture not only intra-beat but inter-beat relationships. Additionally, we observed that larger patch sizes reduce the total number of patches, leading to less stable outcomes. Figure 9 compares ECG reconstructions generated by our proposed ViT-based method and the 1D patch-based Transformer. The results demonstrate that our method produces ECG signals significantly closer to the ground truth. This improvement is attributed to the ability of 2D patches to capture variations in ECG cycles, an essential capability that the 1D sequence-based Transformer lacks.